OpenStudy (anonymous):

Why is 15t (t-5) the answer to 15t^2-75t?

OpenStudy (anonymous):

becoz on opening bracket we multiply the terms

OpenStudy (anonymous):

a(b-a)=ab-a^2

OpenStudy (anonymous):

here 15t.t-15t.5

OpenStudy (anonymous):

that is 15t^2-75t

OpenStudy (anonymous):

@niine Did you understand danish's answer?

OpenStudy (anonymous):

When you multiply exponents you add them \[t \times t = t^2\] it works with normal numbers to, keep in mind everything is to the one, so look. \[2^1 \times 2^1 = 2^{1+1} = 2^2 = 4\] And for your question you have to distribute the 15t over everything in the brackets, as so.